OpenStudy (unicornsession):

1) What is the coefficient of x in the division (18x^3 + 12x^2 - 3x) + 6x^2 3 2 -0.5 -3 2) Divide (2x^3 - x^2 - 24x + 12) divided by (2x-1) x^2 - 1 x^2 - 12 2x - 1 x^2 - 1/2x - 12 3) Simplify 2a+2 over a^2-1 2 over a - 1 2a +2 over a^2 - 1 -2 a+1 over a-1

OpenStudy (anonymous):

need help?

OpenStudy (unicornsession):

yes

OpenStudy (anonymous):

okies im here to help

OpenStudy (anonymous):

aswner is : - Simplifying 18x3 + -12x2 + 6x = 0 Reorder the terms: 6x + -12x2 + 18x3 = 0 Solving 6x + -12x2 + 18x3 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '6x'. 6x(1 + -2x + 3x2) = 0 Ignore the factor 6.

OpenStudy (anonymous):

that helped?

OpenStudy (unicornsession):

not really. i now know 2) is x^2 - 12

OpenStudy (anonymous):

ok ill try again

OpenStudy (unicornsession):

3) 2/a-1

OpenStudy (anonymous):

Simplifying 18x3 + -12x2 + 6x = 0 Reorder the terms: 6x + -12x2 + 18x3 = 0 Solving 6x + -12x2 + 18x3 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '6x'. 6x(1 + -2x + 3x2) = 0 Ignore the factor 6. Subproblem 1 Set the factor 'x' equal to zero and attempt to solve: Simplifying x = 0 Solving x = 0 Move all terms containing x to the left, all other terms to the right. Simplifying x = 0 Subproblem 2 Set the factor '(1 + -2x + 3x2)' equal to zero and attempt to solve: Simplifying 1 + -2x + 3x2 = 0 Solving 1 + -2x + 3x2 = 0 Begin completing the square. Divide all terms by 3 the coefficient of the squared term: Divide each side by '3'. 0.3333333333 + -0.6666666667x + x2 = 0 Move the constant term to the right: Add '-0.3333333333' to each side of the equation. 0.3333333333 + -0.6666666667x + -0.3333333333 + x2 = 0 + -0.3333333333 Reorder the terms: 0.3333333333 + -0.3333333333 + -0.6666666667x + x2 = 0 + -0.3333333333 Combine like terms: 0.3333333333 + -0.3333333333 = 0.0000000000 0.0000000000 + -0.6666666667x + x2 = 0 + -0.3333333333 -0.6666666667x + x2 = 0 + -0.3333333333 Combine like terms: 0 + -0.3333333333 = -0.3333333333 -0.6666666667x + x2 = -0.3333333333 The x term is -0.6666666667x. Take half its coefficient (-0.3333333334). Square it (0.1111111112) and add it to both sides. Add '0.1111111112' to each side of the equation. -0.6666666667x + 0.1111111112 + x2 = -0.3333333333 + 0.1111111112 Reorder the terms: 0.1111111112 + -0.6666666667x + x2 = -0.3333333333 + 0.1111111112 Combine like terms: -0.3333333333 + 0.1111111112 = -0.2222222221 0.1111111112 + -0.6666666667x + x2 = -0.2222222221 Factor a perfect square on the left side: (x + -0.3333333334)(x + -0.3333333334) = -0.2222222221 Can't calculate square root of the right side. The solution to this equation could not be determined. This subproblem is being ignored because a solution could not be determined. Solution x = {0}

OpenStudy (unicornsession):

that's not even one of the choices

OpenStudy (anonymous):

no im just sayin its a example Example problem like urs so now u canc how they do it then u do same way they did it ^